# DBudgetKV: Dynamic Budget in KV Cache Compression for Ensuring Optimal Performance

### v2

@@ 5 Conclusion
In this study, we introduce a novel KV cache compression objective designed to approximate the full-cache performance, independent of specific inputs, while optimizing resource utilization through targeted KV cache pruning. Our approach, termed DBudgetKV, employs a straightforward yet effective two-step process for each layer of LLMs. Tokens are initially ranked solely based on their positions, followed by the eviction of the least significant tokens as determined by the norm value of the reduced attention matrix. Comprehensive experiments conducted across diverse datasets, encompassing a variety of tasks and context lengths, demonstrate that DBudgetKV achieves nearly lossless compression, while with notable space and time reduction, successfully fulfilling our objective for greater practical values.
Appendix A Datasets Used in Experiments
Math & Science
Commonsense Reasoning (CR)
Single Document QA (Single-Doc QA)
This task assesses the model’s reading comprehension skills when dealing with a single, extended document. NarrativeQA [31] is a dataset designed to evaluate model’s ability to comprehend and answer questions based on narrative texts, focusing on understanding stories and their underlying themes. Qasper [32] is a dataset aimed at assessing model’s capability to extract and answer questions from academic papers, emphasizing understanding complex scientific information. We employ F1 Score as the metric for above two datasets.
Multi-Document QA (Multi-Doc QA)
This task evaluates the model’s reading comprehension capabilities across multiple extended documents. 2WikiMultiHopQA (2WKMQA) [33] is a dataset designed to test model’s ability to perform multi-hop reasoning and answer complex questions using information from multiple Wikipedia articles. MuSiQue [34] evaluates model’s skill in integrating and reasoning over information from multiple sources to answer comprehensive questions accurately. We leverage F1 Score as the metric for above two datasets.
Summarization
This task examines the model’s ability to comprehend and summarize lengthy documents. QMSum [35] is a dataset for evaluating model’s ability to generate concise summaries of meeting transcripts, focusing on capturing the key points from multi-party discussions. Multi-News (M-News) [36] is a dataset that challenges models to create coherent summaries by synthesizing information from multiple news articles on the same topic. We use Rouge-L as the metric for above two datasets.
Few-Shot Learning (FSL)
This task assesses the model’s few-shot learning capabilities. TriviaQA [37] is a dataset designed to assess model’s ability to retrieve and answer questions based on large collections of trivia, emphasizing comprehension and factual recall. We use F1 Score as the metric.
Code
This task evaluates the model’s ability to complete and generate code. LCC [38] is a dataset focused on evaluating models’ ability to understand and generate code by considering extended code contexts, enhancing the ability to reason over complex programming structures. We use Edit Sim as the metric.
Appendix B Ablation
@@
Appendix C Freeze the first two layers of LLM
In this section, we demonstrate the necessity of using DBudgetKV to compress the KV cache starting from the 2nd layer of LLMs. We apply DBudgetKV on Llama2-7B-Chat and conduct experiments on GSM8K and CoQA. We first perform a case study, followed by a comparison of the effects of not freezing layers of the model on its performance.
As shown in Table 9, when the threshold is set to 1% and no layers are frozen, the outputs for two examples on GSM8K and CoQA are incorrect and lack logical coherence. However, the budget exceeds 90% in both cases. By observing the actual budget of each layer, we can see that for these two examples, the budgets for layer 0 and 1 are relatively low, while the budgets from layer 2 to layer 30 are the same and very high, with the last layer being low again. We analyze that since the model, in the first two layers, is not yet able to identify truly important tokens, the attention distribution is relatively uniform. This may lead to the model discarding truly important tokens during eviction, resulting in subsequent generations being unable to access this information, ultimately causing the output to fail.
Table 8: Performance on GSM8K and CoQA using DBudgetKV with different frozen layers.
Datasets Layer None 0 0,1 0,1,2 0,1,2,3 0,1,31
GSM8K Oursk=1 0.014 0.250 0.264 0.252 0.264 0.252
Budget 93.1% 97.0% 95.4% 97.5% 97.7% 98.2%
CoQA Oursk=1 1.47 52.80 53.58 53.58 53.32 53.46
Budget 92.7% 94.7% 95.5% 98.3% 98.4% 99.2%
Based on the above case study, we attempt to freeze certain layers of the model and explore the optimal freezing configuration. We continue to test the results of freezing different layers. The resuls are shown in Table 8. From the table, we can observe that freezing the first two layers achieves a balance between model performance and budget. Moreover, freezing the 31st layer does not significantly enhance the model’s performance and instead leads to an increase in budget. Consequently, DBudgetKV ultimately opts to begin KV cache compression from the 2nd layer of the model.
Appendix D Full Algorithm of DBudgetKV
  Input: Prompt, Threshold $t$
  Output: Compressed KV Cache
  Create Empty List $K_{c},V_{c}$
  for Transformer Layer $L_{i}$ in LLM do
     $Q^{i},K^{i},V^{i}\leftarrow L_{i}(\text{Prompt})$
     $R^{i}\leftarrow$ Postion-Based Importance Rank
     $F_{b}^{i}\leftarrow\operatorname{Frobenius}(A_{last}^{i})$
     $A_{last}^{i}\leftarrow\operatorname{Square}(A_{last}^{i})$
     Reorder $A_{last}^{i}$ by Rank $R$
     $A_{cumsum}^{i}\leftarrow\operatorname{Cumsum}(A_{last}^{i})$
     $A_{cumsum}^{i}\leftarrow\operatorname{Sqrt}(A_{cumsum}^{i})$
     $A_{ratio}^{i}\leftarrow(F_{b}^{i}-A_{cumsum}^{i})/F_{b}^{i}$
     $K_{c}^{i}\leftarrow\operatorname{Compress}K^{i}\text{by}R^{i}[I:]$
     $V_{c}^{i}\leftarrow\operatorname{Compress}V^{i}\text{by}R^{i}[I:]$
     Append $K_{c}^{i},V_{c}^{i}$ to $K_{c},V_{c}$
  end for
  return  $K_{c},V_{c}$
Algorithm 1 DBudgetKV
Table 9: Case Study.
GSM8K
Input
A robe takes 2 bolts of blue fiber a-
nd half that much white fiber. How
many bolts in total does it take?
Budget
Layer 0: 67.71
Layer 1: 82.29
Layer 2$\sim$30: 95.83
Layer 31: 37.50
Avg.: 93.90
Output
I have determined by answering the
answer to the format of bolts bolts b-
olts…(repeat)
Ground-Truth 3
CoQA
Input
You are given a story and a question.
Answer the question as concisely as
you can…Question: What color was
Cotton?
Budget
Layer 0: 42.14
Layer 1: 42.54
Layer 2$\sim$30: 99.19
Layer 31: 79.64
Avg.: 95.03
Output
Question: Question: What is the que-
stion: What is the question:
Groud-Truth White
Appendix E Limitations
The main limitation of DBudgetKV stems from the gap between its current compression budget and the true optimal budget. This can be seen in Table 2 that for certain scenarios, e.g. QMSum with Mistral-7B, DBudgetKV reaches 84.3% budget while 50% budget is also viable with no performance degradation. We regard this gap as room for improvement in future work, facilitating more aggressive KV cache compression while ensuring the full-cache performance.
Another limitation of DBudgetKV is that though it demonstrates almost lossless compression, we rely on empirical experiments and there is no hard guarantee on the full-performance constraint. In Table 2, while both Llama3-8B and Qwen2.5-7B even surpass the full-cache performance, DBudgetKV with Mistral-7B shows trivial degradation of 1.5%. More robust methods could be developed as future work to further strengthen this constraint.
Appendix F Broader Impacts
Our work has made contributions on both a new objective in this line of research, as well as a new method that successfully fulfills this objective. Since all current KV cache pruning methods require a pre-defined KV budget threshold, they either have inconsistent and unstable performance across different input lengths, domains and tasks, or they need careful tuning of their budgets per dataset to ensure the full-cache performance. Our proposed method evaluated on 13 diverse datasets, including short and long-context QA, coding and math, demonstrates its effectiveness on addressing this limitation, which automatically adjusts the budget to reach full-cache performance regardless of the inputs.

### v3

@@ 4 Conclusion
In this study, we introduce a new KV cache compression objective that lifts the threshold dependency, so to achieve input-insensitive pruning for robust inference performance. Towards this objective, we propose a novel method, termed ReFreeKV, which employs a straightforward yet effective two-stage KV cache pruning process. Comprehensive experiments conducted across diverse datasets, encompassing a variety of tasks and context lengths, demonstrate that ReFreeKV achieves nearly lossless compression without input-specific tuning or thresholds, while with notable memory reduction.
Limitations
The main limitation of ReFreeKV stems from the gap between its current compression budget and the true optimal budget. This can be seen in Table 2 that for certain scenarios, e.g. QMSum with Mistral-7B, ReFreeKV reaches 84.3% budget while 50% budget is also viable with no performance degradation. We regard this gap as room for improvement in future work, facilitating more aggressive KV cache compression while ensuring the full-cache objective.
Another limitation of ReFreeKV is that though it demonstrates almost lossless compression, we rely on empirical experiments, and there is no hard guarantee on the degree of degradation. In Table 2, while both Llama3-8B and Qwen2.5-7B even surpass the full-cache performance, ReFreeKV with Mistral-7B shows trivial degradation of 1.5%. More robust methods could be developed as future work to further strengthen the robustness.
Appendix A Full Algorithm of ReFreeKV
Algorithm 1 ReFreeKV
Input: Prompt, Threshold $t$
Output: Compressed KV Cache
 Create Empty List $K_{c},V_{c}$
for Transformer Layer $L_{i}$ in LLM do
  $Q^{i},K^{i},V^{i}\leftarrow L_{i}(\text{Prompt})$
  $R^{i}\leftarrow$ Postion-Based Importance Rank
  $F_{b}^{i}\leftarrow\operatorname{Frobenius}(A_{last}^{i})$
  $A_{last}^{i}\leftarrow\operatorname{Square}(A_{last}^{i})$
  Reorder $A_{last}^{i}$ by Rank $R$
  $A_{cumsum}^{i}\leftarrow\operatorname{Cumsum}(A_{last}^{i})$
  $A_{cumsum}^{i}\leftarrow\operatorname{Sqrt}(A_{cumsum}^{i})$
  $A_{ratio}^{i}\leftarrow(F_{b}^{i}-A_{cumsum}^{i})/F_{b}^{i}$
  $K_{c}^{i}\leftarrow\operatorname{Compress}K^{i}\text{by}R^{i}[I:]$
  $V_{c}^{i}\leftarrow\operatorname{Compress}V^{i}\text{by}R^{i}[I:]$
  Append $K_{c}^{i},V_{c}^{i}$ to $K_{c},V_{c}$
end for
return $K_{c},V_{c}$
Appendix B Retain the first two layers of LLM
In this section, we demonstrate the design choice of ReFreeKV to compress the KV cache starting from the third layer of LLMs. We apply ReFreeKV with Llama2-7B-Chat and conduct experiments on GSM8K and CoQA. We first perform a case study, followed by a comparison of the effects of not retaining full KV cache of the first two layers.
As shown in Table 6, when the threshold is set to 1% and no layers are frozen, the outputs for two examples on GSM8K and CoQA are incorrect and lack logical coherence. However, the budget exceeds 90% in both cases. By observing the actual budget of each layer, we can see that for these two examples, the budgets for layer 1 and 2 are relatively low, while the budgets from layer 3 to layer 31 are the same and very high, with the last layer being low again. We analyze that since the model, in the first two layers, is not yet able to identify truly important tokens, the attention distribution is relatively uniform. This may lead to the model discarding truly important tokens during eviction, resulting in subsequent generations being unable to access this information, ultimately causing the output to fail.
Based on the above case study, we attempt to retain certain layers of the model and explore the optimal configuration. We continue to test the results of retain different layers. The results are shown in Table 7. We can observe that freezing the first two layers achieves a balance between model performance and budget. Moreover, retaining the 31st layer does not significantly enhance the model’s performance and instead leads to an increase in budget. Consequently, ReFreeKV ultimately opts to begin KV cache compression from the 3rd layer of the model.
GSM8K
Input
A robe takes 2 bolts of blue fiber a-
nd half that much white fiber. How
many bolts in total does it take?
Budget
Layer 1: 67.71
Layer 2: 82.29
Layer $\sim3$1: 95.83
Layer 32: 37.50
Avg.: 93.90
Output
I have determined by answering the
answer to the format of bolts bolts b-
olts…(repeat)
Ground-Truth 3
CoQA
Input
You are given a story and a question.
Answer the question as concisely as
you can…Question: What color was
Cotton?
Budget
Layer 1: 42.14
Layer 2: 42.54
Layer $\sim3$1: 99.19
Layer 32: 79.64
Avg.: 95.03
Output
Question: Question: What is the que-
stion: What is the question:
Groud-Truth White
Table 6: Case Study.
Datasets Layer None 0 0,1 0,1,2 0,1,2,3 0,1,31
GSM8K Oursk=1 0.014 0.250 0.264 0.252 0.264 0.252
Budget 93.1% 97.0% 95.4% 97.5% 97.7% 98.2%
CoQA Oursk=1 1.47 52.80 53.58 53.58 53.32 53.46
Budget 92.7% 94.7% 95.5% 98.3% 98.4% 99.2%
Table 7: Performance of Llama2-7B-Chat on GSM8K and CoQA using ReFreeKV with different frozen layers.
Appendix C Main Results with 50% Budget
GSM8K
GPQA
TheoQA
ThQA
CoQA
NrtvQA
Qasper
2WkMQA
Musique
QMSum
M-News
TriviaQA
Lcc
Llama3 Full 75.28 29.02 21.29 25.59 52.74 24.06 43.91 35.33 14.77 22.27 27.37 70.26 19.16 100%
Mistral Full 33.36 29.24 6.83 20.82 39.68 28.74 37.80 33.87 22.88 22.19 22.94 86.87 16.08 100%
Qwen2.5 Full 88.02 31.70 29.85 24.55 61.43 20.81 43.17 47.15 30.70 23.64 24.24 87.64 2.44 100%
Table 8: Performance of ReFreeKV and its comparison with five KV cache pruning models on 13 datasets. The cache size is set to 50%.
We show peformance comparison between ReFreeKV and different eviction-based methods with 50% cache size. The results are show in Table 8. From the table, we can observe the same conclusions with Section 3.
@@ Appendix C Main Results with 50% Budget
BS=1 BS=2 BS=8 BS=16
[Input, Output] [4K, 8K] [8K, 16K] [512, 4K] [4K, 8K] [512,512] [4K, 4K] [512, 512] [2K, 2K]
Latency (s/100tokens)
HF Accelerate 3.61 5.29 3.50 5.26 3.48 12.23 4.71 13.01
Ours 3.19 4.08 2.84 3.60 3.06 10.42 4.25 10.64
Throughput (token/s)
HF Accelerate 27.71 18.89 57.08 38.01 230.00 65.41 339.43 122.98
ReFreeKV 31.35 24.51 70.30 55.56 261.83 76.81 376.89 150.40
Budget 73.3% 55.9% 62.0% 34.0% 76.7% 65.3% 82.8% 78.2%
Table 10: Performance comparison across various batch sizes and sequence lengths. We report latency (s/100 tokens, lower is better) and throughput (tokens/s, higher is better). In the header, ’k’ denotes 1024 tokens (e.g., 4K = 4096).
Appendix D Datasets Used in Experiments
Math & Science
Commonsense Reasoning (CR)
Single Document QA (Single-Doc QA)
This task assesses the model’s reading comprehension skills when dealing with a single, extended document. NarrativeQA DBLP:journals/tacl/KociskySBDHMG18 is a dataset designed to evaluate model’s ability to comprehend and answer questions based on narrative texts, focusing on understanding stories and their underlying themes. Qasper DBLP:conf/naacl/DasigiLBCSG21 is a dataset aimed at assessing model’s capability to extract and answer questions from academic papers, emphasizing understanding complex scientific information. We employ F1 Score as the metric for above two datasets.
Multi-Document QA (Multi-Doc QA)
This task evaluates the model’s reading comprehension capabilities across multiple extended documents. 2WikiMultiHopQA (2WKMQA) DBLP:conf/coling/HoNSA20 is a dataset designed to test model’s ability to perform multi-hop reasoning and answer complex questions using information from multiple Wikipedia articles. MuSiQue DBLP:journals/tacl/TrivediBKS22 evaluates model’s skill in integrating and reasoning over information from multiple sources to answer comprehensive questions accurately. We leverage F1 Score as the metric for above two datasets.
Summarization
This task examines the model’s ability to comprehend and summarize lengthy documents. QMSum DBLP:conf/naacl/ZhongYYZMJACLQR21 is a dataset for evaluating model’s ability to generate concise summaries of meeting transcripts, focusing on capturing the key points from multi-party discussions. Multi-News (M-News) DBLP:conf/acl/FabbriLSLR19 is a dataset that challenges models to create coherent summaries by synthesizing information from multiple news articles on the same topic. We use Rouge-L as the metric for above two datasets.
Few-Shot Learning (FSL)
This task assesses the model’s few-shot learning capabilities. TriviaQA DBLP:conf/acl/JoshiCWZ17 is a dataset designed to assess model’s ability to retrieve and answer questions based on large collections of trivia, emphasizing comprehension and factual recall. We use F1 Score as the metric.
Code
This task evaluates the model’s ability to complete and generate code. LCC DBLP:conf/icml/GuoXD0M23 is a dataset focused on evaluating models’ ability to understand and generate code by considering extended code contexts, enhancing the ability to reason over complex programming structures. We use Edit Sim as the metric.
Appendix E Ablation
@@
Appendix F Analysis of Latency and Throughput
We perform following experiments for a comprehensive analysis of the latency and throughput. We follow the setup in FastGen using inputs of NarrativeQA, and conduct end-to-end latency comparison experiments on a single 80GB A100 GPU, using the standard HuggingFace (HF) Accelerate library as the baseline. The results are shown in Table 10.
GSM8K GPQA CoQA NrtvQA QMSum TriviaQA
Full 75.28 29.02 52.74 24.06 22.27 70.26
Oursk=1 76.50 30.13 52.86 23.44 21.95 64.24
Budget 93.2% 86.7% 81.5% 48.7% 15.0% 41.0%
SLM0.9 72.78 28.79 52.83 23.75 22.48 69.97
SLMF0.9 76.42 29.24 52.89 22.95 22.47 69.10
SLM0.5 3.41 6.03 44.68 20.94 21.80 67.79
SLMF0.5 4.55 3.13 44.59 21.31 21.88 68.04
SLM0.2 3.21 3.35 40.11 20.12 21.11 66.17
SLMF0.2 1.21 1.12 38.44 19.52 20.61 65.69
Table 11: Performance comparison between ReFreeKV and StreamingLLM with the first two layers frozen (SLMF) under difference KV cache budgets ($0.9$, $0.5$ and $0.2$).
Methods GSM8K CoQA NQA Musique QMSum
Llama3-70B-Instruct
Full 89.69 60.36 27.15 29.31 22.52
Oursk=1 89.76 60.38 26.94 28.88 22.30
Budget 93.3% 77.2% 58.7% 60.8% 52.9%
Qwen2.5-32B-Instruct
Full 91.36 58.03 24.78 40.04 22.84
Ours 91.81 57.29 22.65 40.54 22.52
Budget 91.6% 85.0% 68.2% 74.4% 76.8%
Qwen2.5-72B-Instruct
Full 90.22 54.14 24.36 42.13 23.93
Ours 90.30 54.19 24.10 41.70 23.31
Budget 90.9% 87.6% 63.4% 65.0% 68.6%
Table 12: Performance of LLMs with scales of 70B, 34B, and 72B across five datasets using ReFreeKV.
From the table, we can observe that ReFreeKV consistently outperforms the baseline in both latency and throughput. Moreover, ReFreeKV holds its advantage when the batch size increases, and its throughput remains robust. It is noteworthy that Table 4 in Section 4.2 also demonstrates the trivial overhead of ReFreeKV due to its simple matrix computation during its two-stage pruning process.
Appendix G Comparison with StreamingLLM
Appendix H Related Work
KV Cache Compression
To mitigate the large memory footprint of the KV cache during LLM inference, a prominent line of work has focused on pruning, which selectively discards less important tokens. Methods like Scissorhands DBLP:conf/nips/LiuDLWXXKS23 and SnapKV DBLP:conf/nips/LiHYVLYCLC24 retain tokens based on high attention scores. Others employ strategies based on token recency and historical importance, such as H2O DBLP:conf/nips/Zhang00CZC0TRBW23 or StreamingLLM DBLP:conf/iclr/XiaoTCHL24. FastGen DBLP:conf/iclr/Ge0LZ0024 further refines this by adapting retention strategies on a per-head basis.
A common limitation of these pruning methods is their reliance on a pre-defined memory budget, which can lead to inconsistent performance across different domains and datasets. In contrast, our work lifts this constraint, aiming for near full-cache performance with a dynamically sized budget without requiring task-specific tuning.
Beyond pruning, other compression paradigms have been explored. These include quantization-based methods, which reduce the numerical precision of cached values DBLP:journals/corr/abs-2405-06219; DBLP:conf/nips/HooperKMMSKG24; DBLP:journals/corr/abs-2503-18773; DBLP:journals/corr/abs-2402-12065; DBLP:journals/corr/abs-2502-04420, and low-rank approximation techniques that apply matrix decomposition to compress the KV cache DBLP:conf/iclr/ChangLLCHWHCAW25; DBLP:journals/corr/abs-2405-04434.
Attention Patterns
Our approach is grounded in research on the structural patterns of attention mechanisms. Studies have shown that attention heads often specialize, with some exhibiting predictable, position-biased behaviors such as consistently focusing on adjacent tokens DBLP:journals/corr/abs-1911-02150. This indicates a degree of redundancy in their learned patterns. Furthermore, the importance of individual heads can vary significantly across tasks DBLP:journals/corr/abs-2404-07965, suggesting opportunities for dynamic optimization.